OpenStudy (laddiusmaximus):

find a and b if the equation of the tangent line to the graph of the function f(x)=x^(4)+ax^(2)+b at x=1 is y=-4x+3 this still has me pretty confused

OpenStudy (anonymous):

f'(x)=4x^3+2ax f'(1)=-4=>4+2a=-4=>a=-4 now x=1=>y=-4x+3=-1 substitute in f(x)=>-1=1-4+b=>b=2

OpenStudy (laddiusmaximus):

kind of lost me there.

OpenStudy (anonymous):

the slope of y = -4x + 3 is -4 so, -4 = f'(1) = 4 + 2a -4 = 4 + 2a -4 -4 = 2a -8 = 2a -4 = a no from original equaation F(x) = x^(4)+ax^(2)+b -1 = 1^4 + -4(1)^2 + b -1 = 1 + -4 + b -1 - 1 + 4 = b b = 2

OpenStudy (dumbcow):

the derivative of a function at any point gives the slope of the tangent line the tangent line will be of the form: y = mx +b m = slope = f'(x) y = f(x)